The current price of an item is $7.36, due to a price change increase of 45.17%. Determine original price.
Let x = Original Price.
Formula used: Original Price = Current Price / (1 + % Change)
x = 7.36 / (1 + 45.17%) / 100
x = 7.36 / 1.4517
x = 5.0699 = 5.07
Original price is less. Where am I correct ?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


$5.07 is the correct answer.

Check:
45.17% of $5.07 = 0.4517*5.07 = 2.290119 which rounds to 2.29
The price has gone up by $2.29
So we go from 5.07+2.29 = 7.36 which confirms the answer.

Or as a slight shortcut:
1.4517*5.07 = 7.360119 which rounds to 7.36

Answer by ikleyn(52784)

.

You are correct everywhere.